A voltaic cell consists of a Pb/Pb2+ half-cell and a Cu/Cu2+ half-cell at 25 ∘C. The initial concentrations of Pb2+ and Cu2+ are 5.30×10−2 M and 1.70 M, respectively.

Part A- What is the initial cell potential? Express your answer using two significant figures.

Part B- What is the cell potential when the concentration of Cu2+ has fallen to 0.230 M ? Express your answer using two significant figures.

Part C- What are the concentrations of Pb2+ and Cu2+ when the cell potential falls to 0.36 V? Enter your answers numerically separated by a comma. Express your answers using two significant figures.
Answer #1

Solution : - Given datas- pbat a 5-30x10 ²M co²t = 1.70 M. PART: A pool phe case, (São *02) Il company (1 70)/eve) Epost P -0Fcell = 0.5low Ecell = 0.3lov / Ran: Eqution - - 1 mol cult produce 1 mol pb²+ Let, se a change in come of both 30/.0.37 .= 0& concentration of cost 0.230m of chang 17 -0.230 = 1047 et concentration of pb I 0.053 t1.47 1.523 L Go - Ecelleme 0.0591 -